OpenStudy (anonymous):

n ^{2}logn and n^{2} which has larger growth rate??

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Yes, remember \(\left(a^n\right)^r=a^{nr}\).

OpenStudy (anonymous):

n^{2}logn has larger growth rate than n^2.

OpenStudy (anonymous):

Well, I'll give a better answer. Comparing \(n^{2\ln n},n^2\), we can say \(n^{2\ln n}=\left(n^2\right)^{\ln n}>n^2\mid\forall\left(n>e\right)\).

OpenStudy (anonymous):

thank u